OpenStudy (anonymous):

integration of sec²(x/2)-1

OpenStudy (luigi0210):

Any limits? Or is it just an indefinite integral?

OpenStudy (anonymous):

no just an indefinite integral

OpenStudy (luigi0210):

@whpalmer4 @Loser66 c:

OpenStudy (loser66):

\[\int (sec^2(x/2)-1dx=\int sec^2(x/2)dx -\int dx\] for the second term, it's easy, right? it's = x +C for the first term , just let u = x/2 du = 1/2 dx or dx = 2du so, it becomes \[\int 2sec^2u du= 2tanu = 2 tan (x/2)\] DAT SIT!! don't forget combine 2 terms

OpenStudy (anonymous):

thank u

OpenStudy (loser66):

oh, I see you there, but didn't pay attention, I thought you are one of the helper. hehehe. ok, np